OpenStudy (anonymous):

is 6/24 equivalent to 1/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because all you do is Simplify \[\frac{ 6 }{ 24 }\]

OpenStudy (e.mccormick):

For this sort of thing you need to factor the fractions to see if you can cancel things and have them end up the same. In this particular case you can divide the top and bottom of the first fraction by 6 because you have a 1 on top of the second. 6/6 = 1, so if 24/6=4 then they would be the same.